if father is 4 times as old as his old as his son, after 24 years the father is twice as old as his son. find the current age of father and son?
Found 2 solutions by jorel1380, ewatrrr:
Answer by jorel1380(3719)

x=4y
x+24=2(y+24)
4y+24=2y+48
2y=24
y=12
x=48
The father is 48. His son is 12.

Answer by ewatrrr(24785)

 
Hi
father is 4 times as old as his old as his son
Let x and 4x represent the son and father's ages NOW repectively
Question states***after 24 yrs...
4x+24 = 2(x+24)
Solving for x
2x = 24
x = 12, the son's age now. His father is 48 now
CHECKING our Answer***after 24 yrs...
72 = 2*36 = 72